OpenStudy (anonymous):

prove that D[arccotx]=-(1/(1+x^2)

OpenStudy (anonymous):

whats captial D mean?

OpenStudy (anonymous):

Derivative

OpenStudy (anonymous):

It's a shorter way of saying dy/dx essentially

OpenStudy (anonymous):

ah ok, ty

OpenStudy (anonymous):

can you use the definition of a derivative to prove it?

OpenStudy (anonymous):

If I knew the properties of arccot, then yeah....but I don't

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

1. Write down y=arccot(x). 2. Draw a picture of a right triangle with base x, height 1, and angle y (between the base and hypotenuse). The hypotenuse will therefore be sqrt(1+x^2). This triangle represents the relationship between x and y and will be important later on. 3. Take the cotangent of both sides to give x=cot(y). 4. Rewrite cot(y) as cos(y)/sin(y). 5. Implicitly differentiate the function. I hope you know how to do this. If you don't, the basic rule is to differentiate both x and y terms as you normally would, except all terms with a y in them are multiplied by y' (where y' is dy/dx). Here's how it's done with implicit differentiation and the quotient rule: x = cos(y)/sin(y) 1 = (sin(y)*(-sin(y)*y') - cos(y)*cos(y)*y')/sin^2(y) 1 = (-sin^2(y)*y' - cos^2(y)*y')/sin^2(y) 1 = y'*(-sin^2(y) - cos^2(y))/sin^2(y) 1 = -y'/sin^2(y) (This is thanks to the identity sin^2(y)+cos^2(y)=1) 6. Now get the derivative term-- the y' term-- on one side of the equation: y' = -sin^2(y) 7. Look at your drawing of the triangle. To get a proper value for the sin, you must divide the length of each side by sqrt(1-x^2) so that the hypotenuse is length 1. Now the base of the triangle is x/sqrt(1+x^2) and the height is 1/sqrt(1+x^2). 8. Remember that the sin of a function is the length of a side opposite the angle on a right triangle with hypotenuse 1. In other words, the sin of y is 1/sqrt(1+x^2) 9. Substitute that into your equation y' = -sin^2(y) to give y' = -1/(1+x^2)